# Time-inversion of spatiotemporal beam dynamics using uncertainty-aware latent evolution reversal

### v1

@@ Abstract
Charged particle dynamics under the influence of electromagnetic fields is a challenging spatiotemporal problem. Many high performance physics-based simulators for predicting behavior in a charged particle beam are computationally expensive, limiting their utility for solving inverse problems online. The problem of estimating upstream six-dimensional phase space given downstream measurements of charged particles in an accelerator is an inverse problem of growing importance. This paper introduces a reverse Latent Evolution Model (rLEM) designed for temporal inversion of forward beam dynamics. In this two-step self-supervised deep learning framework, we utilize a Conditional Variational Autoencoder (CVAE) to project 6D phase space projections of a charged particle beam into a lower-dimensional latent distribution. Subsequently, we autoregressively learn the inverse temporal dynamics in the latent space using a Long Short-Term Memory (LSTM) network. The coupled CVAE-LSTM framework can predict 6D phase space projections across all upstream accelerating sections based on single or multiple downstream phase space measurements as inputs. The proposed model also captures the aleatoric uncertainty of the high-dimensional input data within the latent space. This uncertainty, which reflects potential uncertain measurements at a given module, is propagated through the LSTM to estimate uncertainty bounds for all upstream predictions, demonstrating the robustness of the LSTM against in-distribution variations in the input data.
keywords:
reverse latent evolution , CVAE , autoregressive LSTM , spatiotemporal dynamics , accelerator physics , aleatoric uncertainty
journal: XYZ
[…]
[inst1]Applied Electrodynamics Group (AOT-AE), Los Alamos National Laboratory, USA
1 Introduction
Particle accelerator are complex, time-varying and high-dimensional systems, governed by hundreds to thousands of radio-frequency (RF) cavity and magnet parameters. Charged particles beam dynamics in accelerators is a spatiotemporal dynamical phenomenon with a temporally evolving six-dimensional phase space $(x,y,z,p_{x},p_{y},p_{z})$. Solving such dynamical systems can be computationally demanding due to the intricate temporal evolution intertwined with detailed spatial features.
@@ 1 Introduction
In recent years, machine learning (ML) has shown promising abilities in solving complex problems in physical sciences [1, 2, 3, 4, 5, 6]. However, majority of the research is inclined towards learning spatial or temporal dynamics with limited emphasis on learning full spatiotemporal dynamics. Some of the techniques for solving spatiotemporal dynamical problems are Graph Neural Networks (GNNs) [7], Convolutional Long Short-Term Memory (ConvLSTM) [8],Deep Convolutional Generative Adversarial Networks (DCGAN) [9], three-dimensional Convolutional Neural Networks (3DCNN) [10].
@@
Latent evolution models have recently gained traction for solving challenging spatiotemporal dynamics problems. In these machine learning models, the learning process is decomposed into two subproblems for training, which are later integrated to make predictions [11]. First, a dimensionality reducer captures spatial correlations by projecting high-dimensional images into a lower-dimensional latent space. Then, another model learns the temporal correlations within this latent space. Latent evolution models become computationally efficient by working in lower dimensional space. The encoder-decoder framework is more effective at extracting representative features compared to an encoder network. Consecutively, this improves the learning of temporal correlations using recurrent neural networks. In [12], a convolutional autoencoder (AE) combined with an adaptive feedback is proposed for time-varying beam dynamics in particle accelerators. In [13], principal component analysis and LSTM are employed for learning phase-field-based microstructure evolution. In [11, 14, 15, 16], AE and LSTM are utilized for addressing fluid flow problems [11, 14, 15, 16]. In [17], $\beta$-variational autoencoders (VAE) and transformers are used for reduced-order modelling of fluid flow. In [18, 19], conditional VAE with autoregressive LSTMs are proposed for charged particle beam dynamics in accelerators.
@@
Deep learning has been used regularly to solve forward problems like beam diagnostics in particle accelerators. In [20], a first demonstration of adaptive ML was performed combining adaptive feedback with deep neural networks for automatic shaping of the $(z,E)$ longitudinal phase space of the LCLS Free Electron Laser’s electron beam. In [21] a virtual diagnostic was developed. In [22], polynomial neural networks are developed with symplectic regularization to represent Taylor maps of particle dynamics. Deep Lie map networks were developed to identify magnetic field errors based on beam position monitor measurements in synchrotrons [23]. In [24], virtual diagnostics were developed for accelerator injector tuning. In [25], deep reinforcement learning was used to optimize the design of a superconducting radio-frequency gun. In [26], a method was developed for predicting the transverse emittance of space charge-dominated beams and demonstrated. In [27], ML methods have been combined with multilinear regression to create virtual time of arrival and beam energy diagnostics at HiRES. In [28], uncertainty quantification for virtual diagnostics in particle accelerators is discussed. In [29], a convolutional autoencoder compresses the data into lower dimension and used to relate the phase-space information of 4D phase space distribution to the beam loss at four different locations.
@@
Deep learning is also implemented to solve inverse problems of calibration, fault analysis, tuning and optimization. In [30], Bayesian optimization was used for calibration of intense x-ray FEL pulses. In [31], interpretable ML-based models were developed for predicting breakdowns in high-gradient cavity data. In [32], conditional siamese neural network (CSNN) to predict errant beam under different system configurations. In [33], DL is used to classify superconducting radio-frequency (SRF) cavity faults. In Ref. [34], a genetic algorithm enhanced by ML is used for multi-objective optimization for the dynamic aperture of a storage ring. In [35], time series forecasting methods and their applications to particle accelerators are discussed. In [36], a forward surrogate model of beam bulk parameters at different locations along the accelerator is used for GA-based multi-disciplinary optimization.
@@
1.1 Summary of main results
@@
In general, we propose a reverse latent evolution model (rLEM) designed to predict the upstream 6D phase space from downstream phase space measurements. This two-step self-supervised framework operates as follows: first, a conditional VAE learns the low-dimensional latent distribution of the 15 unique projections of the charged particle beam’s phase space. Next, a Long Short-Term Memory (LSTM)-based recurrent neural network is trained autoregressively to capture the reverse the forward temporal dynamics in the latent space. The coupled network is used to predict phase space projections in upstream accelerating sections based on limited phase space measurements from downstream sections. Our model effectively captures aleatoric uncertainty from the image space and translates it into the latent space. It also demonstrates robustness against in-distribution variations in the input data. While the proposed technique can address general spatiotemporal problems, we have implemented it specifically for a charged particle beam in the linear accelerator at the Los Alamos Neutron Science Center (LANSCE) at Los Alamos National Laboratory (LANL).
2 Methods
2.1 Charged particles beam dynamics
In a linear accelerator, the dynamics of charged particles evolve temporally in a six-dimensional phase space, consisting of three position components and three momentum components for each particle: $(x,y,z,p_{x},p_{y},p_{z})$ with $z$ typically chosen along the accelerator axis [52]. An ensemble of billions of particles is generally described by a time-varying density function $\rho(x,y,z,x^{\prime},y^{\prime},E,t)$ where $(x^{\prime},y^{\prime})=(p_{x}/p_{z},p_{y}/p_{z})$ are the angles at which the particles are moving relative to the accelerator axis. The dynamics of a beam’s phase space density function are governed by the relativistic Vlasov equation
$\frac{\partial\rho}{\partial t}$ $=$ $-\mathbf{v}\cdot\nabla_{\mathbf{x}}\rho+\frac{\partial\mathbf{p}}{\partial t}\cdot\nabla_{\mathbf{p}}\rho,$ (1)
$\frac{\partial\mathbf{p}}{\partial t}$ $=$ $q\left(\mathbf{E}(x,y,z,t)+\mathbf{v}\times\mathbf{B}(x,y,z,t)\right),$ (2)
with $\mathbf{v}$ and $\mathbf{p}$ representing velocity and relativistic momentum:
$\mathbf{v}$ $=$ $\left(\frac{dx}{dt},\frac{dy}{dt},\frac{dz}{dt}\right),\quad\mathbf{p}=(p_{x},p_{y},p_{z})=\gamma m\mathbf{v},$
$\gamma$ $=$ $\frac{1}{\sqrt{1-v^{2}/c^{2}}},\quad v=|\mathbf{v}|,$ (3)
where $c$ is the speed of light. The electromagnetic fields, $\mathbf{E}$ and $\mathbf{B}$ each have external and beam-based sources. The beam-based sources are the electromagnetic fields of the beams themselves, as defined by Maxwell’s equations, which act back on the particles via space charge forces and distort the 6D phase space. The external sources are the electromagnetic devices such as resonant accelerating structures and magnets throughout a particle accelerator which accelerate and focus the beam. At large accelerators such as LANSCE, where the beam is very intense, the powerful self fields of the beam distort and filament the beam’s phase space via nonlinear space charge forces resulting in highly complex phase space distributions.
Measuring a beam’s 6D phase space is essential for precisely controlling and optimizing beam parameters, but it has proven to be nearly impossible so far. The first iterative method for measuring a beam’s 6D phase space was based on 5.7 million individual beams and required 32 hours 6 years ago, the method can now achieve the same measurement in approximately 18 hours [53]. Fast, single-shot measurements, which provide only 2D projections like (x,y) or (z,E) are typically destructive. These methods are mainly used for electron beams and face significant challenges with high-intensity proton beams, which can damage measurement devices. Additional limitations include resolution constraints with short beam pulses, brief run times, and distribution shifts over time due to changes in accelerator components and initial beam distribution.
Particle accelerators face severe limitations in detailed non-invasive phase space diagnostics. Physics-based simulations such as [54, 55, 56, 57] are commonly employed for beam diagnostics in linear accelerators. Physics based simulators can be used for beam diagnostics but available simulators are computationally expensive and require domain knowledge and expert supervision [54, 55, 56, 57]. This limitation restricts their applicability for uncertainty analysis and inverse problems [43]. Therefore, developing computationally efficient and robust probabilistic surrogate models is critical for accurately predicting beam dynamics.
2.2 Multi-particle tracking simulations
@@
In this study, we have used High Performance Simulator (HPSim), an advanced, open-source tool developed at LANL which enables rapid, online simulations for multiple-particle beam dynamics [56]. It solves the Vlasov-Maxwell equations to compute the impact of external accelerating and focusing forces on the charged particle beam, including space charge forces within the beam. We simulate beam dynamics of  1M macroparticles in the LANSCE linear accelerator at Los Alamos National Laboratory [58].
The LANSCE accelerator provides high-intensity $H^{+}$ and $H^{-}$ particle beams up to 800 MeV for various scientific experiments. A layout of the LANSCE linear accelerator is shown in Fig. 1. The LANSCE accelerator includes H+ and H- injectors, each featuring an ion source and a Cockcroft–Walton accelerator, to produce a 750 KeV beam. A beam transport system focuses, guides, and bunches the beams. Then, the 201.25-MHz Drift Tube linac (4 modules) accelerates the beam to 100 MEV, followed by a 805 MHz coupled-cavity linac (additional 44 modules) to accelerate the beams to 800 MeV. The beam is then used for applications in areas such as ultracold neutron source (UCN) studies, proton radiography (PRAD) for dynamic material imaging, and high-energy physics experiments. Additionally, beams can be directed to the proton storage ring (PSR) for creating highly intense beam bunches, used in the weapons neutron research (WNR) facility or the Lujan center for neutron scattering [59].
HPSim contains a realistic representation of the true beam used at LANSCE, able to simulate the beam as it passes through 48 modules of the Drift Tube Linac DTL and the Coupled Cavity Linac CCL. Each module includes two RF settings: one for controlling the phase of the electromagnetic wave responsible for beam acceleration, and another for adjusting the beam’s amplitude. To collect data from HPSim, first, the RF set points (amplitude and phase) for the first four modules are randomly sampled from a uniform distribution keeping the rest of the set points of 44 modules at a mean value. The sampled 96 dimensional vector is entered into the HPSim to simulate the beam dynamics across the length of the accelerator. We take the raw 6D HPSim output and project the particles to 15 unique projections ($x-p_{x}$, $x-y$, $x-p_{y}$, $x-z$, $x-p_{z}$, $y-p_{x}$, $y-p_{y}$, $y-z$, $y-p_{z}$, $z-p_{x}$, $z-p_{y}$, $z-p_{z}$ ($E-\phi$), $p_{x}-p_{y}$, $p_{x}-p_{z}$, and $p_{z}-p_{y}$) at each of the 48 modules.
We have performed 1400 and 500 different simulations in the HPSim to collect the training and the test set, respectively. 3 out of 15 phase space projections ($x-y$, $E-\phi$, and $x^{\prime}-y^{\prime}$) for one of the simulation is plotted in Fig. 1. The plots are transformed to a logarithmic scale to compress the wide range of pixel intensity levels into a more compact range, enhancing visualization clarity. The projections evolve as they interact with RF electromagnetic fields while traversing various modules. Greater variations in projections are observed in the earlier modules, with the variations gradually decreasing as the beam progresses through the later modules.
Figure 1: A layout of the LANSCE linear accelerator and the related scientific areas like IPF, UCN/PRAD, WNR, Lujan, Area A, PSR. HPSim gives realistic representation of the true beam used at LANSCE. Three out of fifteen projections ($x-y$, $E-\phi$, $x^{\prime}-y^{\prime}$) of the 6D phase space of charged particle beams are shown at different modules.
2.3 Reverse latent evolution model
The beam dynamics described by Eqs. 1, 2 encodes evolution of a charged particle beam’s phase space density $\rho$ in the forward time direction. For notational simplicity we refer to $\rho$ at different times as $\{\rho_{1},\rho_{2},\dots,\rho_{M}\}$, where $M$ is the final accelerating module. Since we are working with the phase space of charged particles denoted as $X$ rather than phase space density $\rho$, we adopt notations using $X$ instead of $\rho$. The LANSCE accelerator has 48 different accelerating modules, therefore, the temporal discretization or time steps becomes 48 ($X_{1},X_{2},\dots,X_{48}$).

### v2

@@ Abstract
Charged particle dynamics under the influence of electromagnetic fields is a challenging spatiotemporal problem. Many high performance physics-based simulators for predicting behavior in a charged particle beam are computationally expensive, limiting their utility for solving inverse problems online. The problem of estimating upstream six-dimensional phase space given downstream measurements of charged particles in an accelerator is an inverse problem of growing importance. This paper introduces a reverse Latent Evolution Model (RLE) designed for temporal inversion of forward beam dynamics. In this two-step self-supervised deep learning framework, we utilize a Conditional Variational Autoencoder (CVAE) to project 6D phase space projections of a charged particle beam into a lower-dimensional latent distribution. Subsequently, we autoregressively learn the inverse temporal dynamics in the latent space using a Long Short-Term Memory (LSTM) network. The coupled CVAE-LSTM framework can predict 6D phase space projections across all upstream accelerating sections based on single or multiple downstream phase space measurements as inputs. The proposed model also captures the aleatoric uncertainty of the high-dimensional input data within the latent space. This uncertainty, which reflects potential uncertain measurements at a given module, is propagated through the LSTM to estimate uncertainty bounds for all upstream predictions, demonstrating the robustness of the LSTM to random perturbations in the input.
preprint: APS/123-QED
I Introduction
Particle accelerators are complex high-dimensional systems with hundreds to thousands of components which include radio-frequency (RF) resonant cavities that accelerator the charged particle beams and magnets that focus and steer the beams. The control and optimization of accelerators is not only challenging due to their complexity, but also because both the accelerator components and the initial conditions of their beams drift with time. Furthermore, for any given accelerator condition, the charged particle beam dynamics evolve in a six-dimensional position and momentum phase space $(x,y,z,p_{x},p_{y},p_{z})$ and experience complex collective effects such as space charge forces in which the electromagnetic fields of the beam’s own particles perturb its shape and energy in a highly nonlinear way. To numerically estimate the spatiotemporal evolution of a beam in an accelerator is computationally demanding because of the large length scales involved. For example, while the individual particles in a charged particle bunch which are experiencing collective space charge effects can be bounded by a sphere with a radius of millimeters the bunch also traverses a kilometer long machine as it is accelerated by the external electromagnetic fields of RF cavities and magnets.
@@ I Introduction
In recent years, machine learning (ML) has shown promising abilities in solving complex problems in physical sciences [1, 2, 3, 4, 5, 6]. However, most of the research has focused on either spatial or temporal dynamics with limited emphasis on spatiotemporal dynamics. Some of the techniques for solving spatiotemporal dynamical problems are Graph Neural Networks (GNNs) [7], Convolutional Long Short-Term Memory (ConvLSTM) [8],Deep Convolutional Generative Adversarial Networks (DCGAN) [9], three-dimensional Convolutional Neural Networks (3DCNN) [10].
@@
Latent evolution models have recently gained traction for solving challenging spatiotemporal dynamics problems. In these machine learning models, the learning process is decomposed into two subproblems for training, which are later integrated to make predictions [11]. First, a dimensionality reducer captures spatial correlations by projecting high-dimensional images into a lower-dimensional latent space. Then, another model learns the temporal correlations within this latent space. Latent evolution models become computationally efficient by working in lower dimensional space. The encoder-decoder framework is effective at extracting representative features, which improves the learning of temporal correlations using recurrent neural networks. In [12], adaptive feedback acting on the latent embedding of a convolutional autoencoder (AE) is proposed for time-varying beam dynamics in particle accelerators. In [13], principal component analysis and LSTM are employed for learning phase-field-based microstructure evolution. In [11, 14, 15, 16], AE and LSTM are utilized for addressing fluid flow problems [11, 14, 15, 16]. In [17], $\beta$-variational autoencoders (VAE) and transformers are used for reduced-order modelling of fluid flow. In [18, 19], conditional VAE with autoregressive LSTMs are proposed for charged particle beam dynamics in accelerators.
@@
Deep learning methods are used to solve forward and inverse problems in beam physics and diagnostics. In [20], a model-independent feedback coupled with a neural network is demonstrated for automatic control of longitudinal phase space (LPS) of electron beams. In [21], a ML-based virtual diagnostic was developed to predict LPS distribution. In [22], a physics-based ML approach is discussed where polynomial neural networks with symplectic regularization is proposed to represent Taylor maps of particle dynamics. Deep Lie map networks were developed to identify magnetic field errors based on beam position monitor measurements in synchrotrons [23]. In [24], an encoder-decoder based virtual diagnostics were developed to model the longitudinal phase-space-diagnostic beamline at the photoinector of the European XFEL. In [25], a fully connected neural network was developed to evaluate measured phase advance scan data of space charge-dominated beams. In [26], ML methods have been combined with multilinear regression to create virtual time of arrival and beam energy diagnostics at HiRES. In [27], a convolutional autoencoder compresses the data into lower dimension and used to relate the phase-space information of 4D phase space distribution to the beam loss at four different locations.
@@
Deep learning is also implemented to solve inverse problems of calibration, fault analysis, tuning and optimization. Bayesian optimization (BO) is popularly used for accelerator tuning [28, 29, 30, 31, 32]. Deep reinforcement learning is another widely used method for control and tuning of different accelerator facilities [33, 34, 35]. Inverse problem of fault analysis in accelerators using machine learning is a critical problem of growing interest. In [36], explainable ML-based models using SHapley Additive exPlanation (SHAP) values were developed for predicting breakdowns in high-gradient cavity. In [37], supervised machine learning based conditional siamese neural network (CSNN) and conditional variational auto encoder (CVAE) models to predict errant beam pulses under different system configurations. In [38], manual feature engineering based ML for classifying superconducting radio-frequency (SRF) cavity faults. In Ref. [39, 40, 41], genetic algorithm based multi-objective optimization is demonstrated for accelerator problems. In [42], time series forecasting methods and their applications to particle accelerators are discussed.
@@
I.1 Summary of main results
In this paper, we propose a reverse latent evolution model (RLE) designed to predict 2D projections of the upstream 6D phase space from downstream phase space measurements. This two-step self-supervised framework operates as follows: first, a conditional VAE learns the low-dimensional latent distribution of the 15 unique projections of the charged particle beam’s phase space. Next, a Long Short-Term Memory (LSTM)-based recurrent neural network is trained autoregressively to capture the reverse the forward temporal dynamics in the latent space. The coupled network is used to predict phase space projections in upstream accelerating sections based on limited phase space measurements from downstream sections. Our model effectively captures aleatoric uncertainty from the image space and translates it into the latent space. It also demonstrates robustness against in-distribution variations in the input data. While the proposed technique can address general spatiotemporal problems, we have implemented it specifically for a charged particle beam in the linear accelerator at the Los Alamos Neutron Science Center (LANSCE) at Los Alamos National Laboratory (LANL).
@@ I.1 Summary of main results
Our current research is inspired by previous work [18], where a latent evolution model (LEM) is used to solve the forward problem of forecasting the spatiotemporal dynamics of charged particles. However, the inverse problem of estimating upstream states given downstream measurement is more complex problem. In the literature [43, 44, 45], separate models are often trained to address this inverse problem. In a LEM, spatial and temporal dynamics are learned independently, enhancing its effectiveness for inverse problem-solving. Contrary to [18], our objective is to invert the spatiotemporal dynamics of charged particles. We achieve this by training the LSTM in the latent space in a reverse autoregressive setting while keeping the spatial learner (CVAE) intact. The main contributions of this research include (a) solving the inverse problem of estimating upstream states from downstream measurements using a RLEM, and (b) studying uncertainty by (i) capturing aleatoric uncertainty in the input space within the latent space, and (ii) propagating aleatoric uncertainty in the latent space for enhanced robustness.
II Methods
II.1 Charged particles beam dynamics
Accurate measurement of a beam’s six-dimensional (6D) phase space is one of the most significant challenges in accelerator physics. While such measurements are essential for precise control and optimization of beam parameters, they remain exceptionally difficult to perform. The first iterative method for reconstructing a beam’s 6D phase space, developed six years ago, required 32 hours and over 5.7 million individual beam measurements [59]. Although advancements have reduced this time to approximately 18 hours, the approach is still impractical for real-time applications. Fast, single-shot measurement techniques provide limited 2D projections, such as (x,y) and (z,E), but these are typically destructive and unsuitable for high-intensity proton beams due to potential damage to diagnostic devices. Further challenges include limited resolution for short beam pulses, restricted operation times, and temporal shifts in the beam distribution caused by variations in accelerator components or initial conditions. These constraints underscore the need for robust theoretical and numerical frameworks to analyze beam dynamics and predict phase space evolution.
The beam dynamics is governed by the interaction between charged particles and electromagnetic fields. At the core of this interaction lies the Vlasov-Maxwell equations, a set of coupled equations describing the self-consistent evolution of charged particle systems in phase space [60]. The state of the system in phase space is represented by both the position $(x,y,z)$ and momentum $(p_{x},p_{y},p_{z})$ of the particles.
The relativistic Vlasov equation governs the evolution of the particle distribution function $f(\mathbf{x},\mathbf{p},t)$, encapsulating the density of particles in phase space:
$\frac{\partial f}{\partial t}+\mathbf{v}\cdot\nabla_{\mathbf{x}}f+\frac{\mathbf{F}}{m}\cdot\nabla_{\mathbf{p}}f=0,$
where the force $\mathbf{F}=q(\mathbf{E}+\mathbf{v}\times\mathbf{B})$ is determined by the electromagnetic fields $\mathbf{E}$ and $\mathbf{B}$. Simultaneously, Maxwell’s equations govern the evolution of these fields:
$\nabla\cdot\mathbf{E}=\frac{\rho}{\epsilon_{0}},\quad\nabla\cdot\mathbf{B}=0,$
×$\nabla\times\mathbf{E}=-\frac{\partial\mathbf{B}}{\partial t},\quad\nabla\times\mathbf{B}=\mu_{0}\mathbf{J}+\mu_{0}\epsilon_{0}\frac{\partial\mathbf{E}}{\partial t}.$
Here, the charge density $\rho=q\int f\,d^{3}\mathbf{p}$ and current density $\mathbf{J}=q\int\mathbf{v}f\,d^{3}\mathbf{p}$ are derived from the particle distribution function. Together, the Vlasov and Maxwell equations form a self-consistent framework, where the particle distribution evolves under the influence of electromagnetic fields, and the fields themselves are shaped by the particles’ charge and current densities.
Electromagnetic fields in accelerators are beam-based sources, which are self-generated fields acting on particles via space charge forces and distorting the 6D phase space, and external sources, such as resonant accelerating structures and magnets, which accelerate, guide, and focus the beam. In high-intensity accelerators, such as the Los Alamos Neutron Science Center (LANSCE), the strong self-fields of the beam result in pronounced nonlinear effects. Nonlinear space charge forces further distort and filament the beam’s phase space, creating highly complex and nonuniform distributions. These intricate dynamics pose significant analytical and computational challenges for accurate modeling and simulation.
II.1.1 Dataset Generation and Description
@@
High Performance Simulator (HPSim) is an open-sourced code developed at LANL which enables rapid, online simulations for multiple-particle beam dynamics [61]. It solves the Vlasov-Maxwell equations to compute the impact of external accelerating and focusing forces on the charged particle beam, including space charge forces within the beam. The software is designed to replicate the accelerator and, therefore, provides a realistic representation of the true beam used at the Los Alamos Neutron Science Center (LANSCE).
We simulate beam dynamics of 1M macroparticles in the LANSCE linear accelerator at Los Alamos National Laboratory. The beam’s behavior in each module is governed by two radio-frequency (RF) parameters: amplitude and phase set-points. The accelerator consists of 48 modules through which the beam propagates, including 4 modules of a 201.25 MHz drift tube linac (DTL) and 44 modules of an 805 MHz coupled cavity linac (CCL).
The dataset was generated using HPSim by varying the RF set points (8 parameters in total) which were randomly sampled from a uniform distribution, while the RF set points for the remaining 44 modules (88 parameters) were kept constant. The resulting 96-dimensional RF set point vector was used as input to HPSim to simulate beam dynamics across the accelerator’s 48 modules. For each simulation, the raw 6D phase space output of HPSim was projected onto 15 unique 2D planes, including $x-p_{x}$, $x-y$, $x-p_{y}$, $x-z$, $x-p_{z}$, $y-p_{x}$, $y-p_{y}$, $y-z$, $y-p_{z}$, $z-p_{x}$, $z-p_{y}$, $z-p_{z}$, $E-\phi$, $p_{x}-p_{y}$, $p_{x}-p_{z}$, and $p_{z}-p_{y}$, for every module.
In total, 1400 simulations were performed to generate the training dataset, and an additional 700 simulations were conducted for testing. The data output from each single simulation has a size of [48, 15, 256, 256] where at each of the 48 modules there are 15 projections, each at a $256\times 256$ pixel resolution. The module indices (m) and projections (X) were normalized to the range [0, 1]. The dataset has been made publicly available as an open-source resource on Zenodo [62].
A high level overview of LANSCE and RF module locations, as well as a sample of the dataset are presented in Fig. 1. Three of the 15 projections ($x-y$, $E-\phi$, $x^{\prime}-y^{\prime}$) are shown at various locations along LANSCE. The figure highlights the complex, nonlinear, and multi-scale spatiotemporal evolution observed in the projections of the 6D beam phase space.
Figure 1: Three 2D projections ($x-y$,$E-\phi$,$x^{\prime}-y^{\prime}$) out of 15 projections of 6d phase space of charged particle beam in the LANSCE linear accelerator. Accelerating modules - 1 to 4 are 201 MHz drift tube linac (DTL) and 5 to 48 are 805 MHz coupled cavity linac (CCL). The beam serves various scientific areas like isotope production facility (IPF), ultra-cold neutrons (UCN), proton radiography (PRAD), weapons neutron research (WNR), proton storage ring (PSR).
II.2 Reverse latent evolution model (RLEM)
The forward discretized spatiotemporal beam dynamics (described above) can be written as $X_{t}=H(X_{1},X_{2},\dots,X_{t},\dots,X_{T-1})$, where $H$ is a unknown nonlinear function that can be learned as $P(X_{T}|X_{1},X_{2},\dots,X_{t},\dots,X_{T-1})$ with a latent evolution model as presented in [18].